Exercise 6-10 (Part Level Submission)

Consider the following independent situations.

Collapse question part

(a)

Henry Finley wishes to become a millionaire. His money market fund has a balance of $149,668 and has a guaranteed interest rate of 8%. How many years must Henry leave that balance in the fund in order to get his desired $1,025,000? (Round answer to 0 decimal places, e.g. 45.)

years
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1,025,000=149,668*(1.08)^n

(1,025,000/149,668)=(1.08)^n

Taking log on both sides;

log (1,025,000/149,668)=n*log 1.08

n=log (1,025,000/149,668)/log 1.08

=25 years(Approx).